Janet invested 10,000, part at 15% and part at 7%. If the total interest at the end of the year is $1,260, how much did she invest @ 15%?
a) 3,000
b) 8,000
c) 7,000
d) 6,000
Show all work

Answer by ptaylor(1894)


Interest(I) equals Principal(P) times Rate(R) times Time(T) or I=PRT
Let x=amount invested at 15%
Then 10,000-x=amount invested at 7%
Interest at 15% =x*0.15*1 or 0.15x
Interest at 7%=(10,000-x)*0.07*1=0.07(10,000-x)
Now we are told that the total interest for the year was $1260 so our equation to solve is:
0.15x+0.07(10,000-x)=1260 get rid of parens
0.15x+700-0.07x=1260 subtract 700 from both sides
0.15x+700-700-0.07x=1260-700 collect like terms
0.08x=560 divide both sides by 0.08
x=$7000----amount invested at 15%
10,000-x=10,000-7000=$3000---------------amount invested at 7%
CK
7000*0.15+3000*0.07=1260
1050+210=1260
1260=1260
Hope this helps---ptaylor